The average mark of candidates in an aptitude test was 128.5 with a standard deviation of 8.2. Three scores extracted from the test are 148,102,152. What is the average to the extracted scores that are extreme values(outliers)
1
Expert's answer
2020-06-02T18:26:18-0400

"Z=\frac{x- \mu}{\sigma}\\
Z_{1}=\frac{148- 128.5}{8.2}=2.38\\
Z_{2}=\frac{102- 128.5}{8.2}=-3.23\\
Z_{3}=\frac{152- 128.5}{8.2}=2.86\\
\text{We can treat any data value with a z-score less}\\
\text{ than -2 or greater than 2 as an outlier}\\
\text{ so all of them are outliers.}\\
\text{the average of the extracted scores that }\\
\text{are extreme values (outliers)}
\\=\frac{148+102+152}{3}=134"